Change the equation 5x + 4y = -4 to slope-intercept form
Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

Change the equation 5x + 4y = -4 to slope-intercept form
------------
Solve for y.
That's all it takes.
-------
5x + 4y = -4
4y = -5x - 4
y = (-5/4)x - 1 is slope-intercept form

Answer by ikleyn(53937)

.

The slope-intercept form is y = ax + b.

You do the rest.